OpenStudy (anonymous):

I'm given f (x) = 2x + 3 and g (x) = x squared + 1 and I need to find (gof) (x) = g (f(x)). How do I go about solving this equation?

zepdrix (zepdrix):

\[\large \color{royalblue}{g(x)=x^2+1}\] \[\large \color{orangered}{f(x)=2x+3}\] Composition of functions. So what we're going to do is, we're going to take the orange function, and plug it into the blue function. Let's think for a moment about what happens when you plug values in for x, into the g function. \[\large \color{royalblue}{g(\color{green}{x})=\color{green}{x}^2+1}\]\[\large \color{royalblue}{g(\color{green}{2})=\color{green}{2}^2+1}\] We replace every x with the given value, right? We're going to be doing the same thing here, but instead of plugging in 2, we'll be plugging in f(x) to replace our x.

OpenStudy (anonymous):

so now it's g(2) = 5

zepdrix (zepdrix):

In the example? Yes good :) That's just an example though.

OpenStudy (anonymous):

oh ok

zepdrix (zepdrix):

\[\large \color{royalblue}{g(\color{green}{x})=\color{green}{x}^2+1}\] So let's plug f(x) in place of our green x. \[\large \color{royalblue}{g(\color{orangered}{f(x)})=\color{orangered}{f(x)}^2+1}\]

zepdrix (zepdrix):

It looks a little messy I know, but understand what we're doing? Instead of plugging a number in for x, we're plugging in a function.

OpenStudy (precal):

nice use of color @zepdrix

OpenStudy (anonymous):

Ya it's starting to make a little sense

zepdrix (zepdrix):

\[\large \color{royalblue}{g(\color{orangered}{f(x)})=\color{orangered}{f(x)}^2+1}\] Let's remember what our f(x) was, we'll be wanting to plug that in for f(x).\[\large \color{orangered}{f(x)=2x+3}\] Plugging this in gives us,\[\large \color{royalblue}{g(\color{orangered}{f(x)})=\color{orangered}{(2x+3)}^2+1}\]

OpenStudy (anonymous):

So, now would it be g (f (x) ) = 4x squared + 9 + 1 ?

zepdrix (zepdrix):

Woops you didn't expand out the square properly. Remember that \(\large (2x+3)^2=(2x+3)(2x+3)\) You should get 4 terms coming out of that multiplication in total - two of which you can combine.

OpenStudy (anonymous):

Oh right...it should be 4x squared + 12x + 9

zepdrix (zepdrix):

Yay good job \c:/

OpenStudy (anonymous):

So, am I done with the problem, or is there still more to do?

zepdrix (zepdrix):

Ah sorry got distracted. You are indeed done. \[\large (g \;o\;f)(x)=g(f(x))=4x^2+12x+10\]

OpenStudy (anonymous):

Alright thanks. I just did another problem which required me to find the equation of a line going through points (-2,3) and (5,6). Would I use the distance formula for this one?

OpenStudy (precal):

no use find the slope first, then use y=mx+b to find b sub your m and b values into y=mx+b

OpenStudy (anonymous):

So, use the midpoint formula?

OpenStudy (precal):

no, y=mx+b you should close out this problem and post your second problem as a new one

OpenStudy (anonymous):

I figured it out